# PromptGuard: Soft Prompt-Guided Unsafe Content Moderation for Text-to-Image Models

### v2

@@ 6 Discussion and Conclusion
Drawing inspiration from the system prompt mechanism in LLMs, our study investigates an innovative content moderation technique that can be highly efficient and lightweight while generating images, termed PromptGuard. It demands no additional models or introduces perturbation during the diffusion denoising process, achieving minimal computational and time overhead. To overcome the lack of a direct system prompt interface in the T2I models, we optimize the safety pseudo-word acting as an implicit system prompt, guiding visual latent away from unsafe regions in the embedding space. Our divide-and-conquer strategy, careful data preparation, and loss function further enhance moderation across varied NSFW categories. Our extensive experiments compare eight state-of-the-art defenses, achieving an optimal unsafe ratio as low as 5.84%. Furthermore, we demonstrate that PromptGuard is 7.8 times more efficient than previous content moderation methods.
Limitation. Despite its strengths, our work is limited by the absence of user involvement in experiments, because we are careful with unsafe content and avoid its exposure to participants due to ethical considerations. Therefore, the statistics of NSFW removal rate is conducted using an open-source NSFW classifier [35], which might result in inherent measurement errors across each experimental outcome. Nevertheless, our manual validation demonstrates these results accurately delineate the relative efficacy of different defenses.
Future Work. This work focuses on the alignment of text-to-image (T2I) models and aims to promote responsible AI practices. We believe that our lightweight moderation can be extended to other generative models, such as text-to-video and image-to-image models, to prevent the generation of NSFW content in these areas.
Supplementary Material
7 Additional Experiment Setup
7.1 Test Benchmark
@@
  • SneakyPrompt: To compensate for the shortcomings of I2P dataset in pornographic data, we use the NSFW dataset curated by [46] as the sexual part of our benchmark.
@@
To apply the I2P dataset to our classification of unsafe categories, we first need to reclassify the data. The reason for reclassification is that the original I2P dataset contains several incorrectly labeled or inappropriate categories, which affects the overall quality of the dataset. Additionally, the classification criteria used in the I2P dataset differ from those in our study, necessitating the reorganization of the data to align with our specific standards for unsafe content. We achieve this by leveraging GPT4-o [4] as a classifier, using the instruction within this box.
7.2 Evaluation Metrics
The additional details of four metrics used to evaluate PromptGuard and other baselines are as follows:
@@ 7.3 Baselines
  • [Model Alignment] SafeGen: we use the official pretrained weights provided in [20] to generate images.
@@ 7.4 Implementation Details
During training, we separately optimize the soft prompt embeddings for each of the four unsafe categories and combine them into a $4\times N$ dimensional token embedding, where $N$ represents the dimensionality of the token embedding space of the CLIP text encoder used in SDv1.4. For inference, the individual embeddings are concatenated and appended to the end of the input prompt token embeddings.
8 Additional Evaluation Results
8.1 Impact of $\lambda$ Across NSFW Categories
Similar to the results and analysis in Section 5.5.1, increasing the value of $\lambda$ encourages $P_{*}$ to lose its ability to generate unsafe images during latent denoising. Figure 6 illustrates the variations in images generated by the model with embeddings trained using different values of $\lambda$.
@@ 8.1 Impact of $\lambda$ Across NSFW Categories
8.2 NSFW Content Moderation
@@ 8.3 Benign Preservation
8.4 Cross-Category Generalization of Individual Soft Prompt Embedding
@@
The results, shown in Table 8, reveal notable differences in generalization across the four unsafe categories. Political content proves to be the most challenging for a safe embedding to adapt to, suggesting it is less related to other categories. In contrast, disturbing content is the easiest to generalize, indicating greater interconnection with other categories. An intriguing observation is that embeddings trained on violent data underperform on violent test data relative to those trained on sexual content. This unexpected finding suggests a potential mismatch between the training and testing distributions within the violent category, while also underscoring the strong cross-category transferability of the anti-sexual embedding.
@@
Table 8: Performance of each individual safe embedding transferred to other unseen NSFW categories.
@@
8.5 Transfer our framework on other T2I models
Stable Diffusion V1.5 The Stable-Diffusion-v1-5 checkpoint was initialized from Stable-Diffusion-v1-2 and fine-tuned for 595k steps at a resolution of 512x512 on the “laion-aesthetics v2 5+” dataset, with 10% dropout of text-conditioning to improve classifier-free guidance. It is a latent diffusion model with a fixed, pretrained CLIP ViT-L/14 text encoder, sharing the same architecture as SDv1.4. Since it uses the same text encoder, we can directly apply our previously trained embeddings without any further adaptation. The test results are shown in Table 9.
@@ 8.5 Transfer our framework on other T2I models
Table 9: Performance of directly applying embeddings trained on SDv1.4 to SDv1.5 for NSFW moderation. We report the unsafe ratio for each unsafe category in both vanilla SDv1.5 and SDv1.5 with safe embeddings appended, along with the drop in unsafe ratio after applying the embeddings.
@@
Stable Diffusion XL Stable Diffusion XL (SDXL) [34] is an enhanced latent diffusion model designed for high-quality text-to-image synthesis. Unlike its predecessor, Stable Diffusion v1.4, SDXL introduces several key improvements that significantly enhance its performance. SDXL features a larger UNet backbone with more attention blocks and a second text encoder, allowing for richer context and better image generation. Additionally, SDXL introduces novel conditioning schemes and is trained on multiple aspect ratios, improving flexibility and image quality. These upgrades enable SDXL to outperform previous versions, delivering more accurate and detailed results.
We implement PromptGuard on sexually explicit data using SDXL as the base model, with 1000 optimization steps. The NSFW moderation performance for different values of the coefficient $\lambda$ is shown in Table 10. We observe that the unsafe ratio for the model protected by PromptGuard, across various $\lambda$ values, shows a notable drop compared to the vanilla SDXL. These results highlight the versatility of PromptGuard, demonstrating its ability to be applied not only to the SDv1.4 model but also to other text-to-image architectures, with consistent effectiveness in enhancing NSFW moderation.
Table 10: Performance of applying PromptGuard with SDXL as base model on sexually explicit unsafe content. We report the unsafe ratio for different $\lambda$, along with the drop in unsafe ratio after applying the embeddings.

### v3

@@ VI Conclusion
Inspired by the system prompt mechanism in large language models (LLMs), we introduce a new content moderation technique for image generation, PromptGuard. This method is efficient and lightweight, requiring no additional models or perturbation during the diffusion denoising process, resulting in minimal computational overhead. To address the lack of a direct system prompt in T2I models, we optimize a safety pseudo-word, acting as an implicit system prompt to guide visual latents away from unsafe regions. Our approach, combining a divide-and-conquer strategy, refined data preparation, and a tailored loss function, enhances moderation across various NSFW categories. Extensive experiments comparing eight state-of-the-art defenses show that PromptGuard reduces the unsafe content ratio to as low as 5.84%. Moreover, PromptGuard is 3.8 times more efficient than previous moderation methods. We also discuss the limitations and future work in VI-C in the appendix.
Appendix
VI-A Additional Experiment Setup
VI-A1 Test Benchmark
@@
  • NSFW-200: To compensate for the shortcomings of I2P dataset in pornographic data, we use the NSFW dataset from [40] for the sexual category.
@@
  • SneakyPrompt: SneakyPrompt [40] is an RL-based attack and we reproduce two variants of it: SneakyPrompt-N with natural words and SneakyPrompt-P with pseudo words to assess the adversarial robustness.
  • MMA-Diffusion: MMA-Diffusion [41] is a dual-modal attack that could bypass safeguards and post-hoc safety checkers using pseudo-words for stealth.
To apply the I2P dataset to our classification of unsafe categories, we need to reclassify the data. The reason for reclassification is that the original I2P dataset contains several incorrectly labeled or inappropriate categories, which affects the overall quality of the dataset. Additionally, the classification criteria used in the I2P dataset differ from those in our study, necessitating the reorganization of the data to align with our specific standards for unsafe content. We achieve this by leveraging GPT4-o [37] as a classifier, using the instruction shown in this box
VI-A2 Evaluation Metrics
The additional details of four metrics used for evaluation are as follows:
@@ VI-A3 Baselines
  • [Model Alignment] SafeGen: We retrain SafeGen using the PromptGuard dataset for Violent, Political, and Disturbing content, creating four separate models for each NSFW category (Sexually Explicit, Violent, Political, and Disturbing) instead of the unified model in the official version. This ensures a fairer comparison by addressing the suboptimal performance of training categories together. While the official SafeGen uses pre-trained weights [44] for sexually explicit content, our version has been retrained for the other three categories to improve moderation accuracy.
@@ VI-A4 Implementation Details
VI-B Additional Evaluation Results
VI-B1 Impact of $\lambda$ Across NSFW Categories
Similar to the results and analysis in V-E1, increasing the value of $\lambda$ encourages $P_{*}$ to lose its ability to generate unsafe images during latent denoising. Figure 6 illustrates the variations in images generated by the model with embeddings trained using different values of $\lambda$.
@@ VI-B1 Impact of $\lambda$ Across NSFW Categories
VI-B2 NSFW Content Moderation
@@ VI-B3 Benign Preservation
VI-B4 Cross-Category Generalization of Individual Soft Prompt Embedding
@@
The results, shown in Table X, reveal notable differences in generalization across the four unsafe categories. Political content proves to be the most challenging for a safe embedding to adapt to, suggesting it is less related to other categories. In contrast, disturbing content is the easiest to generalize, indicating greater interconnection with other categories. An intriguing observation is that embeddings trained on violent data underperform on violent test data relative to those trained on sexual content. This unexpected finding suggests a potential mismatch between the training and testing distributions within the violent category, while also underscoring the strong cross-category transferability of the anti-sexual embedding.
@@
TABLE X: Performance of each individual safe embedding transferred to other unseen NSFW categories.
@@
VI-B5 Exploration on Number of Benign Categories.
Our initial six categories were selected based on the COCO dataset [38]. To further investigate the impact of benign prompt diversity, we introduce two additional categories: Technologies & Electronic Devices and Art & Culture. We then evaluate Benign Preservation performance on sexually explicit training data across different numbers of benign categories. As shown in Table XI, our results indicate that incorporating a more diverse set of benign categories in the training dataset generally improves benign preservation performance, reinforcing the importance of category diversity in maintaining generation quality.
TABLE XI: Benign Preservation of different benign categories.
Number of Benign Categories 4 5 6 7 8
CLIP Score↑ 26.20 26.20 26.28 26.02 26.43
LPIPS Score↓ 0.641 0.638 0.637 0.638 0.636
VI-B6 Transfer our framework on other T2I models
Stable Diffusion V1.5. The Stable-Diffusion-v1-5 checkpoint was initialized from Stable-Diffusion-v1-2 and fine-tuned for 595k steps at a resolution of 512x512 on the “laion-aesthetics v2 5+” dataset, with 10% dropout of text-conditioning to improve classifier-free guidance. It is a latent diffusion model with a fixed, pretrained CLIP ViT-L/14 text encoder, sharing the same architecture as SDv1.4. Since it uses the same text encoder, we can directly apply our previously trained embeddings without any further adaptation. The test results are shown in Table XII.
@@ VI-B6 Transfer our framework on other T2I models
TABLE XII: Performance of directly applying embeddings trained on SDv1.4 to SDv1.5 for NSFW moderation. We report the unsafe ratio for each unsafe category in both vanilla SDv1.5 and SDv1.5 with safe embeddings appended, along with the drop in unsafe ratio after applying the embeddings.
@@
Stable Diffusion XL. Stable Diffusion XL (SDXL) [51] is an enhanced latent diffusion model designed for high-quality text-to-image synthesis. Unlike its predecessor, Stable Diffusion v1.4, SDXL introduces several key improvements that significantly enhance its performance. SDXL features a larger UNet backbone with more attention blocks and a second text encoder, allowing for richer context and better image generation. Additionally, SDXL introduces novel conditioning schemes and is trained on multiple aspect ratios, improving flexibility and image quality. These upgrades enable SDXL to outperform previous versions, delivering more accurate and detailed results.
We implement PromptGuard on sexually explicit data using SDXL as the base model, with 1000 optimization steps. The NSFW moderation performance for different values of the coefficient $\lambda$ is shown in Table XIII. We observe that the unsafe ratio for the model protected by PromptGuard, across various $\lambda$ values, shows a notable drop compared to the vanilla SDXL.
TABLE XIII: Performance of applying PromptGuard with SDXL as base model on sexually explicit unsafe content. We report the unsafe ratio for different $\lambda$, along with the drop in unsafe ratio after applying the embeddings.
@@
DeepFloyd IF. DeepFloyd IF [52] is a novel state-of-the-art open-source text-to-image model with a high degree of photorealism and language understanding. The model is a modular composed of a frozen text encoder and three cascaded pixel diffusion modules. All stages of the model utilize a frozen text encoder based on the T5 transformer [47] to extract text embeddings, which are then fed into a UNet architecture enhanced with cross-attention and attention pooling.
We implement PromptGuard on sexually explicit data using SeepFloyd IF as the base model. The NSFW moderation performance for different values of the coefficient $\lambda$ is shown in Table XIII. We could observe that the unsafe ratio also show a drop under different settings of hyperparameters. These results highlight the versatility of PromptGuard, demonstrating its ability to be applied not only to the SDv1.4 model but also to other text-to-image architectures even beyond CLIP-based latent diffusion models, with consistent effectiveness in enhancing NSFW moderation.
TABLE XIV: Performance of applying PromptGuard with DeepFloyd IF as base model on sexually explicit unsafe content.
coefficient Vanilla DeepFloyd IF 0.1 0.2 0.3 0.4 0.5 0.6 0.7
Unsafe Ratio
45.00 41.00 38.00 25.50 24.00 21.50 36.50 39.00
Unsafe Ratio
Drop (%)↑
/ 4.00 7.00 19.50 21.00 23.50 4.50 6.00
VI-C Additional Discussions
Limitations. A key limitation is the lack of user involvement in our experiments due to ethical concerns about exposing unsafe content. Consequently, NSFW removal rates are measured using an open-source classifier [20], which may introduce minor errors in results. However, manual validation confirms that our results accurately compare the effectiveness of different defenses.
Transferability. In Section VI-B6 we have demonstrated PromptGuard’s ability to transfer to other T2I architectures. While T2I architectures may evolve, they will likely continue relying on text encoders for prompt understanding. Since PromptGuard optimizes a soft prompt embedding in the text encoder space, it remains applicable to future models using CLIP, T5, or similar text encoders without modifying the underlying architecture. Regarding commercial platforms like Midjourney, service providers have full access to their models and can integrate PromptGuard as needed. Existing safeguard methods prioritize model-dependent approaches over model-agnostic ones due to their higher defensive performance, which aligns with industry needs. Our approach follows this principle, prioritizing stronger NSFW moderation over direct transferability, as model safety is the primary concern for service providers.
Future Work. While this study focuses on moderating NSFW content in T2I models, we believe that our lightweight PromptGuard can be extended to other generative models, such as text-to-video and image-to-image models, to prevent the generation of NSFW content in these areas. Moreover, as societal norms and challenges evolve, new unsafe content categories may emerge, necessitating the timely adaptation of moderation frameworks. Furthermore, we hope that future content moderation methods will take a more comprehensive approach by addressing a broader spectrum of unsafe content, ensuring generative models remain aligned with evolving ethical and safety standards.